# Multi-objective Aerial IRS-assisted ISAC Optimization via Generative AI-enhanced Deep Reinforcement Learning

### v1

@@ V-D2 Diffusion Model-based Actor Network
The diffusion model consists of two key processes, i.e., the forward process and the reverse process. In the forward process, the model gradually perturbs the original data by adding noise step by step, transforming the data into a nearly pure noise state. In the reverse process, the model learns the inverse of the noise-adding process, attempting to progressively denoising the noisy data and ultimately recovering the original data. In particular, this progressive denoising process enables the model to capture complex nonlinear relationships and dependencies in the data, thus enabling deep modeling of the data distribution. As such, diffusion model-based DDPG are better able to generate progressively more refined and optimized decisions based on the current environment state during the exploration and exploitation, and enhance its adaptability in complex environments. Next, the detailed mathematical representation of the diffusion model is presented.
Forward Process: Let $x_{0}$ denote the original data, and assume that the forward process consists of $G$ steps, denoted as $\mathcal{G}=\{1,\dots,g,\dots,G\}$. Note that the data at the $g$-th step is expressed as $x_{g}$, which shares the same dimensionality as $x_{0}$. By gradually adding Gaussian noise to $x_{0}$ over $G$ steps, the sequence of data $x_{1},x_{2},\dots,x_{G}$ is obtained. As such, the distribution $q(x_{g}|x_{g-1})$ from $x_{g-1}$ to $x_{g}$ can be expressed as
@@
Figure 2: The architecture of the proposed GDMDDPG for solving CSEMOP.
Reverse Process: The reverse process is a step-by-step denoising procedure aimed at recovering the original data $x_{0}$ from the noisy data $x_{G}$. If the distribution $q(x_{g-1}|x_{g})$ from $x_{g}$ to $x_{g-1}$ is known, $x_{0}$ can be sampled directly. However, statistical estimates of $q(x_{g-1}|x_{g})$ involves complex data distribution calculations, which are difficult to implement in practice. To address this, a parameterized model $p_{\theta}$ is introduced to approximate $q(x_{g-1}|x_{g})$, which can be expressed as
@@
As such, according to Eq. (36) and Eq. (33), the transition distribution $p_{\eta}$ is obtained. Then, the original data $x_{g-1}$ can be sampled from the distribution $p(x_{g})p_{\eta}(x_{g-1}|x_{g})$. Similar to the forward process, the original data $x_{0}$ can be obtained from the distribution as follows:
@@
V-D3 Noise Perturbation Mechanism
@@
2 Initialize experience replay buffer $\mathcal{B}$ and batch size $B$.
@@
9                   Calculate the mean $\kappa_{\eta}$ and variance $\tilde{\beta}_{g}$ based on Eq. (36) and Eq. (33).
@@
15             Sample $B$ transitions from $\mathcal{B}$ according to the probability in Eq. (42), and then update $\eta$ and $\zeta$ according to Eq. (26) and Eq. (44).
@@
V-D4 RPER Sampling Mechanism
@@
Priority-based Sampling: Given that different experiences contribute variably to the training process, priority-based sampling can enhance training efficiency more effectively than simple random uniform sampling. In the RPER sampling mechanism, the priority of each experience sample is quantified by TD error, and experiences are sample from the sampling range defined by Eq. (40). In this case, the priority of the $i$-th experience can be expressed as
@@
$P_{i}=\frac{p_{i}^{\beta_{1}}}{\sum_{j}p_{j}^{\beta_{1}}},\quad i,j\in D_{F_{u}},$ (42)
@@
TABLE I: Simulation Parameters
Parameter Value Parameter Value
Scenario $X_{\min}$ 300 m $Y_{\max}$ 300 m
$V_{\max}$ 20 m/s $L_{0}$ 0.001
$M$ 4 $\eta^{bu}$ 1
$\alpha^{br}$ 2.2 $\alpha^{ru}$ 2.2
$\alpha^{bu}$ 4.6 $\sigma_{n}^{2}$90 dBm
$\sigma_{g}^{2}$90 dBm $L$ 16
$P_{s}$ 79.85 $P_{m}$ 88.63
$U_{r}$ 120 m/s $V_{h}$ 4.03
$d_{0}$ 0.6 $\rho_{a}$ 1.225 kg/m3
$z$ 0.05 $G$ 0.503 m2
DRL $N_{e}$ 4500 $\gamma$ 0.99
$B_{\max}$ 100000 $B$ 128
$F_{\min}$ 3000 $\varepsilon$ 0.005
V-E Main Flow of GDMDDPG Method
Fig. 2 illustrates the architecture of the proposed GDMDDPG method. In this method, the conventional actor network of DDPG is improved by the diffusion model to facilitate the generation of diverse and high-quality actions. Moreover, the noise perturbation mechanism is introduced to further strengthen the exploration of the agent on the action space. Meanwhile, the RPER mechanism improves the algorithm training efficiency by giving higher priority to recent experiences with significant TD-error. The comprehensive steps for implementing the GDMDDPG method are provided in Algorithm 1.
V-E1 Training and Execution
In our considered aerial IRS-assisted ISAC system, the BS possesses sufficient computational capability to undertake the training and deployment of the algorithm. Therefore, the BS is designated as the agent, which implements our proposed GDMDDPG method to make decisions based on the environment states. Moreover, throughout the training process, all generated experiences are retained in a buffer, which is used to update the network parameters. After undergoing sufficient comprehensive training, the BS can adaptively perform real-time operations according to the environment states through the well-trained actor network.
V-E2 Computation Complexity Analysis
The computational complexity of training the GDMDDPG method consists of the following components [26]:
Network Initialization: GDMDDPG method contains an actor network, a critic network, a target actor network, and a target critic network. Assume that the number of parameters for the actor network and the critic network are $|\eta|$ and $|\zeta|$. Moreover, the target network and the main network have the same number of parameters. Therefore, the computation complexity of the network initialization process is $\mathcal{O}(2|\eta|+2|\zeta|)$.
@@ V-E2 Computation Complexity Analysis
As such, the computation complexity of the proposed GDMDDPG method is $\mathcal{O}(2|\eta|+2|\zeta|+N_{e}GT|\eta|+N_{e}TV+N_{e}T(B+2|\eta|+2|\zeta|)$.
VI Simulation Results
In this section, we construct simulations to validate the performance of the proposed GDMDDPG method. First, the parameter settings and the comparison methods are introduced. Subsequently, the experimental results are presented and analyzed in detail.
VI-A Simulation Setup
VI-A1 Environment Details
@@
VI-A2 Network Structure
The proposed GDMDDPG method consists of a critic network and a diffusion model-based actor network. The critic network includes two hidden layers, each comprising 256 neurons and utilizing the ReLU activation function. The input layer of the diffusion model-based actor network handles three types of data, i.e., the reverse step index, the action from the previous step, and the current environment state. To capture temporal dynamics in the diffusion process, sinusoidal positional embeddings are applied. Moreover, the denoising network consists of two hidden layers, each containing 256 neurons with ReLU activation. In addition, more parameters about the proposed GDMDDPG are shown in Table I.
Figure 3: Convergence of the proposed GDMDDPG method with different parameters. (a) Learning rates. (b) Noise schedule functions.
Figure 4: Convergence curves of different methods. (a) Average Reward. (b) Average communication rate. (c) Average echo rate. (d) Average UAV energy consumption.
VI-B Comparison Methods
We evaluate the effectiveness of the proposed method by comparing it with the following methods:
@@ VI-B Comparison Methods
  • Communication-only Method: In this method, the GDMDDPG method focuses only on the communication performance of the system, i.e., it only maximizes the communication rate between the BS and the users.
@@
  • Sensing-only Method: In this method, the GDMDDPG method focuses only on the sensing performance i.e., it only maximizes the echo rate of the target.
@@
Moreover, we also compare our proposed GDMDDPG method with the following State of the art DRL algorithms:
@@
  • DDPG: DDPG algorithm is used to solve the CSEMOP to verify the effectiveness of the proposed improvement factors. Note that DDPG is applied with the noise perturbation mechanism to encourage exploration.
@@
  • TD3: TD3 algorithm can be considered as an enhanced version of DDPG, which employs improvements such as dual Q-network, delayed update, and noise regularization to address the limitations of DDPG.
VI-C Simulation Results
@@
In this part, we analyze the performance of the GDMDDPG method under different setup conditions versus the comparison methods.
VI-C1 Impact of Algorithm Parameters
@@
Fig. LABEL:sub@subfig:Convergence_LR shows the convergence performance of the GDMDDPG method under different learning rates. It is evident that the method performs better in terms of both accuracy and stability when the learning rate is set to $5e^{-4}$. This is because, with a smaller learning rate, the learning process of the agent becomes too slow, often leading to premature convergence to a local optimal. On the contrary, a larger learning rate makes the parameter update step of the neural network become larger, resulting in the algorithm oscillating or even not converging during the training process.
@@
Fig. LABEL:sub@subfig:noise_schedule_function shows the convergence performance of the GDMDDPG method under different noise scheduler functions. The effectiveness of the diffusion model is highly dependent on the noise schedule function, which determines how the noise level changes with the diffusion steps. Specifically, this function influences the ability of the model to learn and generate high-quality samples by controlling the noise variation in the diffusion process. To systematically assess the effect of different noise scheduler functions on the proposed GDMDDPG method, we conducted comparative experiments using three representative scheduling functions., i.e., VP, linear, and cosine. As can be seen, the GDMDDPG method performs better when VP is used as a noise scheduling function, which suggests that it is more helpful in solving the formulated CSEMOP.
Figure 5: The impact of transmit power on different optimization objectives. (a) Average Reward. (b) Average echo rate. (c) Average UAV energy consumption.
VI-C2 Comparison with Other Methods
@@
Fig. 4 shows the convergence curves of all the methods in terms of reward and the three optimization objectives. The results show that the performance of the diffusion model-based DDPG algorithm without the noise perturbation mechanism is significantly degraded, particularly during the initial phase of training, the UAV frequently flies out of the target area. This is due to the fact that DDPG is a deterministic policy algorithm, and the noise perturbation mechanism not only enhances the exploration ability of the DDPG algorithm on the solution space, but also effectively alleviate the issue of outputting boundary actions during the early stage of training. Moreover, it can be found that the TD3 algorithm outperforms the DDPG algorithm. This improvement is attributed to the fact that TD3 mitigates the overestimation issue of Q-values in DDPG by introducing dual Q-networks and enhances training stability through target policy smoothing regularization.
@@
Compared to both DDPG and TD3, the proposed GDMDDPG method achieves superior performance across all three optimization metrics. This performance enhancement stems from three improvement factors proposed in this paper. First, the diffusion model strengthens the agent ability to analyze environment state, thereby improving decision-making quality. Second, noise perturbation mechanism enhances the exploratory capability of the agent within the action space. Finally, the RPER sampling mechanism, which integrates ERE and PER techniques, enables more efficient utilization of experience data. Specifically, it avoids forgetting past experiences while increasing the sampling frequency of recent experiences with high learning value. This experience sampling mechanism further boosts the learning efficiency of the algorithm.
VI-C3 Impact of Maximum Power $P_{max}$
Fig. 5 shows the trends of the three optimization objectives in CSEMOP with the variation of the BS maximum transmit power $P_{max}$. Experimental results demonstrate that as $P_{max}$ increases, the communication rate and target echo rate of DDPG, TD3, and the proposed GDMDDPG methods all show an upward trend. This is because the increase in $P_{max}$ enhances the signal strength, thereby improving both communication and sensing performance. Moreover, the GDMDDPG method achieves the lowest UAV energy consumption, demonstrating its effectiveness in UAV trajectory design.

### v2

@@ 5.4.2 Diffusion Model-based Actor Network
The diffusion model consists of two key processes, i.e., the forward process and the reverse process [44]. In the forward process, the model gradually perturbs the original data by adding noise step by step, transforming the data into a nearly pure noise state. In the reverse process, the model learns the inverse of the noise-adding process, attempting to progressively denoising the noisy data and ultimately recovering the original data. In particular, this progressive denoising process enables the model to capture complex nonlinear relationships and dependencies in the data, thus enabling deep modeling of the data distribution. As such, diffusion model-based DDPG are better able to generate progressively more refined and optimized decisions based on the current environment state during the exploration and exploitation, and enhance its adaptability in complex environments. Next, the detailed mathematical representation of the diffusion model is presented.
Forward Process: We start from original data $x_{0}$ and define a forward noise-injection process composed of $G$ discrete steps, indexed by $\mathcal{G}=\{1,\ldots,g,\ldots,G\}$. At each step $g\in\mathcal{G}$, the perturbed data is represented as $x_{g}$, which shares the same shape as $x_{0}$. Throughout this process, Gaussian noise is gradually introduced, producing a sequence of increasingly noisy samples. This leads to a Markov chain ${x_{0},x_{1},\dots,x_{G}}$, where each transition from $x_{g-1}$ to $x_{g}$ is governed by a conditional distribution $q(x_{g}\mid x_{g-1})$, defined as
@@
Figure 2: The architecture of the proposed GDMDDPG algorithm for solving CSEMOP.
Reverse Process: The reverse process is formulated as a sequential denoising procedure that reconstructs the original sample $x_{0}$ from the noisy input $x_{G}$. Ideally, if the true reverse conditional distribution $p(x_{g-1}\mid x_{g})$ were available for all $g\in\{G,G{-}1,\dots,1\}$, one could recover $x_{0}$ by sampling through the full reverse trajectory. However, exact evaluation of $p(x_{g-1}\mid x_{g})$ is generally intractable due to the complexity of the data distribution. To avoid this issue, a parameterized distribution $p_{\eta}(x_{g-1}\mid x_{g})$ is introduced to approximate the reverse process, which can be defined as
@@
As such, according to Eqs. (36) and (33), the transition distribution $p_{\eta}$ is obtained. Then, the original data $x_{g-1}$ can be sampled from the distribution $p(x_{g})p_{\eta}(x_{g-1}|x_{g})$. Similar to the forward process, the original data $x_{0}$ can be obtained from the distribution as follows:
@@
5.4.3 Noise Perturbation Mechanism
@@
2 Initialize experience replay buffer $\mathcal{B}$, batch size $B$.
@@
9                   Calculate the mean $\kappa_{\eta}$ and variance $\tilde{\beta}_{g}$ based on Eqs. (36) and (33).
@@
16             Sample $B$ transitions from $\mathcal{B}$ according to the probability in Eq. (42), and then update $\eta$ and $\zeta$ according to Eqs. (26) and (44).
@@
5.4.4 RPER Sampling Mechanism
@@
Priority-based Sampling: Given that different experiences contribute variably to the training process, priority-based sampling can enhance training efficiency more effectively than simple random uniform sampling. In the RPER sampling mechanism, the priority of each experience sample is quantified by TD error, and experiences are sampled from the sampling range defined by Eq. (40). In this case, the priority of the $i$-th experience can be expressed as
@@
$P_{i}=\frac{p_{i}^{\beta_{1}}}{\sum_{j\in D_{F_{u}}}p_{j}^{\beta_{1}}},\quad i\in D_{F_{u}},$ (42)
@@
5.5 Main Flow of GDMDDPG algorithm
Fig. 2 illustrates the architecture of the proposed GDMDDPG algorithm. In this algorithm, the conventional actor network of DDPG is improved by the diffusion model to facilitate the generation of diverse and high-quality actions. Moreover, the noise perturbation mechanism is introduced to further strengthen the exploration of the agent on the action space. Meanwhile, the RPER mechanism improves the algorithm training efficiency by giving higher priority to recent experiences with significant TD-error. The comprehensive steps for implementing the GDMDDPG algorithm are provided in Algorithm 1.
5.5.1 Training and Execution
In our considered low-altitude IRS-assisted ISAC system, the BS possesses sufficient computational capability to undertake the training and deployment of the algorithm. Therefore, the BS is designated as the agent, which implements our proposed GDMDDPG algorithm to make decisions based on the environment states. Moreover, throughout the training process, all generated experiences are retained in a buffer, which is used to update the network parameters. After undergoing sufficient comprehensive training, the BS can adaptively perform real-time operations according to the environment states through the well-trained actor network.
5.5.2 Computation Complexity Analysis
The computational complexity of training the GDMDDPG algorithm consists of the following components [29]:
Network Initialization: GDMDDPG algorithm contains an actor network, a critic network, a target actor network, and a target critic network. Assume that the number of parameters for the actor network and the critic network are $|\eta|$ and $|\zeta|$. Moreover, the target network and the main network have the same number of parameters. Therefore, the computation complexity of the network initialization process is $\mathcal{O}(2|\eta|+2|\zeta|)$.
@@ 5.5.2 Computation Complexity Analysis
As such, the computation complexity of the proposed GDMDDPG algorithm is $\mathcal{O}(2|\eta|+2|\zeta|+N_{e}GT|\eta|+N_{e}TV+N_{e}T(B+2|\eta|+2|\zeta|)$.
6 Simulation Results
In this section, we construct simulations to validate the performance of the proposed GDMDDPG algorithm. First, the parameter settings and the comparison methods are introduced. Subsequently, the simulation results are presented and analyzed in detail.
6.1 Simulation Setup
6.1.1 Environment Details
@@
TABLE I: Simulation Parameters
Parameter Value Parameter Value
System Model $X_{\rm\min}$ 300 m $Y_{\rm\max}$ 300 m
$V_{\rm max}$ 30 m/s $L_{0}$ 0.001
$M$ 4 $\eta^{\rm bu}$ 1
$\alpha^{\rm br},\alpha^{\rm ru}$ 2.2 $\alpha^{\rm ru}$ 2.2
$\alpha^{\rm bu}$ 4.6 $\sigma^{2}$90 dBm
$P_{\rm s}$ 79.85 $P_{\rm m}$ 88.63
$U_{\rm r}$ 120 m/s $V_{\rm h}$ 4.03
$d_{0}$ 0.6 $\rho_{\rm a}$ 1.225 kg/m3
$z$ 0.05 $G$ 0.503 m2
DRL $\xi_{1}$ 1000 $p_{\rm o}$ 50
$N_{\rm e}$ 4500 $\gamma$ 0.99
$B_{\max}$ 100000 $B$ 128
$F_{\min}$ 3000 $\varepsilon$ 0.005
6.1.2 Network Structure
The proposed GDMDDPG algorithm consists of a critic network and a diffusion model-based actor network. The critic network includes two hidden layers, each comprising 256 neurons and utilizing the ReLU activation function. The input layer of the diffusion model-based actor network handles three types of data, i.e., the reverse step index, the action from the previous step, and the current environment state. To capture temporal dynamics in the diffusion process, sinusoidal positional embeddings are applied. Moreover, the denoising network consists of two hidden layers, each containing 256 neurons with ReLU activation. In addition, major parameters about the proposed GDMDDPG are shown in Table I.
Figure 3: Optimization objective values obtained by different methods. (a) Average communication rate. (b) Average sensing rate. (c) Average UAV energy consumption.
6.2 Comparison Methods
We evaluate the effectiveness of the proposed algorithm by comparing it with the following methods:
@@ 6.2 Comparison Methods
  • Communication-only Method: In this method, the GDMDDPG algorithm focuses only on the communication performance of the system, i.e., it only maximizes the communication rate between the BS and the users.
@@
  • Sensing-only Method: In this method, the GDMDDPG algorithm focuses only on the sensing performance i.e., it only maximizes the sensing rate of the target.
@@
Moreover, we compare our proposed GDMDDPG algorithm with the following state-of-art DRL algorithms:
@@
  • DDPG: DDPG algorithm is used to solve the CSEMOP to verify the effectiveness of the proposed improvement factors. Note that DDPG is applied with the noise perturbation mechanism to encourage exploration [39].
@@
  • TD3: TD3 algorithm can be considered as an enhanced version of DDPG, which employs improvements such as dual Q-network, delayed update, and noise regularization to address the limitations of DDPG [46].
6.3 Optimization Results
@@
In this part, we analyze the performance of the GDMDDPG algorithm under different setup conditions versus the comparison methods.
6.3.1 Algorithm Performance Evaluation
Fig. 3 presents the optimized objective values achieved by different algorithms. As can be seen, the proposed GDMDDPG algorithm outperforms the conventional DDPG algorithm, TD algorithm, and random method in both communication rate and sensing rate due to its improved environmental understanding and exploration capabilities. Moreover, GDMDDPG demonstrates the best performance in energy consumption optimization, which is crucial for the energy-constrained UAV platform. As such, the superior communication and sensing performance combined with the lowest energy consumption establishes GDMDDPG as a practical and effective algorithm for the considered low-altitude IRS-assisted ISAC system. In addition, we observe that the communication-only and sensing-only methods achieve optimal values in communication rate and sensing rate metrics compared to GDMDDPG, respectively, which confirms that a significant trade-off relationship exists in the optimization objectives in the formulated CSEMOP.
Figure 4: Convergence of the proposed GDMDDPG algorithm with different parameters. (a) Learning rates. (b) Noise schedule functions.
Figure 5: Convergence curves of different methods. (a) Average Reward. (b) Average communication rate. (c) Average sensing rate. (d) Average UAV energy consumption.
6.3.2 Convergence Performance Evaluation
In the following, we analyze the impact of different key hyper-parameters on the convergence of the proposed GDMDDPG algorithm and further compare its convergence with that of different DRL methods.
@@ 6.3.2 Convergence Performance Evaluation
(i) Hyper-parameter Optimization Results: Fig. 4(a) shows the convergence performance of the GDMDDPG algorithm under different learning rates. It is evident that the algorithm performs better in terms of both accuracy and stability when the learning rate is set to $5\times 10^{-4}$. This is because, with a smaller learning rate, the learning process of the agent becomes too slow, often leading to premature convergence to a local optimal. On the contrary, a larger learning rate makes the parameter update step of the neural network larger, resulting in the algorithm oscillating or even not converging during the training process.
@@
Fig. 4(b) shows the convergence performance of the GDMDDPG algorithm under different noise scheduler functions. The effectiveness of the diffusion model is highly dependent on the noise schedule function, which determines how the noise level changes with the diffusion steps. Specifically, this function influences the ability of the model to learn and generate high-quality samples by controlling the noise variation in the diffusion process. To systematically assess the effect of different noise scheduler functions on the proposed GDMDDPG algorithm, we conduct comparative simulations using three representative scheduling functions, i.e., VP, linear, and cosine. As can be seen, the GDMDDPG algorithm performs better when VP is used as a noise scheduling function, which suggests that it is more helpful in solving the formulated CSEMOP.
@@
(ii) Comparison with Other Methods: Fig. 5 shows the convergence curves of all the methods in terms of the reward and three optimization objectives. The results show that the performance of the diffusion model-based DDPG algorithm without the noise perturbation mechanism is significantly degraded, particularly during the initial phase of training, the UAV frequently flies out of the target area. This is due to the fact that DDPG is a deterministic policy algorithm, and the noise perturbation mechanism not only enhances the exploration ability of the DDPG algorithm on the solution space, but also effectively alleviates the issue of outputting boundary actions during the early stage of training. Moreover, it can be found that the TD3 algorithm outperforms the DDPG algorithm. This improvement is attributed to the fact that TD3 mitigates the overestimation issue of Q-values in DDPG by introducing dual Q-networks and enhances training stability through target policy smoothing regularization.
Figure 6: The impact of transmit power on different optimization objectives. (a) Average communication rate. (b) Average sensing rate. (c) Average UAV energy consumption.
Figure 7: The impact of UAV flight height on different optimization objectives. (a) Average communication rate. (b) Average sensing rate. (c) Average UAV energy consumption.
@@
Compared to both DDPG and TD3, the proposed GDMDDPG algorithm achieves superior performance across all three optimization metrics. This performance enhancement stems from three improvement factors proposed in this paper. First, the diffusion model strengthens the agent ability to analyze environment state, thereby improving decision-making quality. Second, noise perturbation mechanism enhances the exploratory capability of the agent within the action space. Finally, the RPER sampling mechanism, which integrates ERE and PER techniques, enables more efficient utilization of experience data. Specifically, it avoids forgetting past experiences while increasing the sampling frequency of recent experiences with high learning value. This experience sampling mechanism further boosts the learning efficiency of the algorithm.
6.4 Impact of System Settings
In the following, we evaluate the performance of different algorithms under different system settings, including the transmit power of the BS, the UAV flight distance, and the number of BS antennas.
6.4.1 Impact of Maximum Transmit Power